OpenStudy (anonymous):

so what if it was log 7 square root of 7

OpenStudy (anonymous):

log 7 + log root 7 log 7 + 1/2 log 7

OpenStudy (anonymous):

m i rite...

OpenStudy (anonymous):

why did you say 1/2 log 7

OpenStudy (anonymous):

coz log root 7 equals log 7^1/2 and by rules of log.. it equals 1/2 log 7

OpenStudy (angela210793):

Is it \[\log_{7} \sqrt{7} \]

OpenStudy (anonymous):

yes

OpenStudy (angela210793):

=log_7,7^(1/2)=1/2*log_7,7=1/2*1=1/2

OpenStudy (angela210793):

\[\log_{7} 7^1/2=1/2*\log_{7} 7=1/2\]

OpenStudy (anonymous):

using the change of base log(base7) sqrt(7) = log(base10) sqrt(7)/ log(base10)7 =1/2